OpenStudy (anonymous):

solve for x: x^2 + 9x = 0

OpenStudy (anonymous):

Factor it for me first.

OpenStudy (anonymous):

Factor out x, set both equations equal to 0, then solve.

OpenStudy (dwelsher):

x=0, x=-9 (x^2+9x=0) (x=0,x+9=0,-----X=0,x=-9

OpenStudy (skullpatrol):

OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (dwelsher):

skullpatrol- thanks for advise

OpenStudy (anonymous):

thankyou :) @dwelsher

OpenStudy (skullpatrol):

@dwelsher np

OpenStudy (skullpatrol):

@xcaseynoel Are you mad at me?

OpenStudy (skullpatrol):

@xcaseynoel If you are I'm sorry, but I'm just trying to prepare you for a test situation.